# Quantum Steenrod operations of symplectic resolutions

### v2

@@ 2.2. Quantum Steenrod operations
Following [Lee23], we implement the incidence constraints via genuine cycles. Fix a cohomology class $b=\mathrm{PD}[Y]\in H^{*}(X;\mathbb{F}_{p})$ that is (a mod $p$ reduction of) Poincaré dual to a locally finite homology cycle $[Y]$ represented by a union $Y=\bigcup Y_{(i)}$ of embedded submanifolds $Y_{(i)}\subseteq X$. Denote the map $Y\to X$ determined by the inclusions $Y_{(i)}\subseteq X$ as $Y\subseteq X$. Also fix $b_{0}\in H^{*}(X)$, $b_{\infty}\in H^{*}_{c}(X)$ Poincaré dual to a locally finite cycle $[Y_{0}]$ and compact cycle $[Y_{\infty}]$, respectively. Note that $b_{\infty}^{\vee}\in H^{*}(X)$ from the Poincaré pairing. Let $Y_{1}=\cdots=Y_{p}=Y$ be copies of the inclusion $Y\to X$.
@@
    For each $w\in E\mathbb{Z}/p$, $\mathcal{Y}_{w}$ is a product of maps $Y_{j}\to X$ representing a cycle homologous to the distinguished inclusion $Y_{j}\subseteq X$; denote the image of this map also by $\mathcal{Y}_{w}$;
@@
(2.15) $\mathcal{M}^{eq}_{A}:=\left\{(u:C\to X,w\in E\mathbb{Z}/p):u_{*}[C]=A,\ (\overline{\partial}_{J}u)_{z}=\nu^{eq}_{w,z,u(z)}\right\}.$
@@
    The evaluation map from moduli space of simple stable maps ¯$\overline{\mathcal{M}}_{A}^{eq}$ is transverse to $\mathcal{Y}^{eq}$, so the compactified moduli space with incidence constraints is also regular.
@@
Suppose 2.4 holds. Fix the unique $i\geq 0$ such that $\dim\mathcal{M}^{eq;i}_{A}\pitchfork\mathcal{Y}^{eq;i}=0$. Then the mod $p$ count of this 00-dimensional moduli space is well-defined.
@@
    $\Sigma_{1}=1$ for $1\in H^{0}(X;\mathbb{F}_{p})$, the unit class.
@@
    $\Sigma_{b}(b_{0})|_{q^{A}=0}=\mathrm{St}(b)\smile b_{0}$, hence $\Sigma_{b}$ deforms the classical Steenrod operations in quantum parameters.
@@
The classical Steenrod operations in fact arise from the action of the symmetric group $S_{p}$ acting on $X^{p}$, not its subgroup $\mathbb{Z}/p\leq S_{p}$ of cyclic permutations. Since $H^{*}_{S_{p}}(\mathrm{pt};\mathbb{F}_{p})\cong\mathbb{F}_{p}[\![t^{p-1},\theta]\!]\hookrightarrow H^{*}_{\mathbb{Z}/p}(\mathrm{pt};\mathbb{F}_{p})\cong\mathbb{F}_{p}{[\![t,\theta]\!]}$, many coefficients the expansion of $\mathrm{St}(b)$ from (2.24) in the equivariant parameters would in fact vanish. However, $\Sigma_{b}$ in general involves all monomials in $t,\theta$ that satisfy the dimension constraints (2.23). ∎
@@
(2.27) $\nabla_{a}\beta=t\partial_{a}\beta-a\ast\beta$
@@ 2.3. $T$-equivariant quantum Steenrod operations
For simplicity, we place the following assumption about the topology of the target manifold $X$, which is indeed satisfied for the examples considered in this paper.
Assumption 2.12.
The singular cohomology $H^{*}(X;\mathbb{F}_{p})$ is concentrated in even degrees, and the nontrivial cohomology classes $b\in H^{*}(X;\mathbb{F}_{p})$ can be represented as Poincaré duals of (locally finite) cycles in $X$ given by unions of embedded submanifolds.
Remark 2.13.
In the presence of a Hamiltonian torus action, the above assumption is not stringent if one is content with considering generic $p$ and pseudocycles in place of genuine cycles; cf. [MT06].
@@
(2.29) $D\rho\circ J^{eq}_{\theta\cdot w^{\prime},x}=J^{eq}_{w^{\prime},\rho(\theta)(x)}\circ D\rho:\ TX_{x}\to TX_{\theta\cdot x}.$
@@ A $(\mathbb{Z}/p\times T)$-equivariant inhomogeneous term is the data of
(2.30) $\nu^{eq}\in C^{\infty}((E\mathbb{Z}/p\times ET)\times C\times X;\mathrm{Hom}^{0,1}(TC,TX))$
@@
(2.31) $\nu^{eq}_{\eta\cdot w,\theta\cdot w^{\prime},z,x}=D\rho(\theta)^{-1}\circ\nu^{eq}_{w,w^{\prime},\tau(\eta)(z),\rho(\theta)(x)}\circ D\tau(\eta):TC_{z}\to TX_{x}.$
@@
Similarly, one can define the following analogue of the incidence cycles from 2.3. Fix a cohomology class $b=\mathrm{PD}[Y]\in H^{*}_{T}(X;\mathbb{F}_{p})$ that is Poincaré dual to a $T$-invariant locally finite homology cycle $[Y]$. Also fix $b_{0}\in H^{*}_{T}(X)$, $b_{\infty}\in H^{*}_{c,T}(X)$ Poincaré dual to a $T$-invariant locally finite cycle $[Y_{0}]$ and $T$-invariant compact cycle $[Y_{\infty}]$, respectively. Again fix the distinguished map $\mathcal{Y}=Y_{0}\times(Y\times\cdots\times Y)\times Y_{\infty}\subseteq X\times X^{p}\times X$; note that the image is $T$-stable.
@@
As the initial choice of $\mathcal{Y}$ is $(\mathbb{Z}/p\times T)$-stable, we can construct the equivariant incidence cycle as a family of perturbations of $\mathcal{Y}$ parametrized over $E{\mathbb{Z}/p}\times ET$; see [Lee23, Lemma 5.5].
@@
(2.33) $\mathcal{M}^{eq}_{A}:=\left\{(u:C\to X,(w,w^{\prime})\in E\mathbb{Z}/p\times ET):u_{*}[C]=A,\ (\overline{\partial}_{J_{w^{\prime}}}u)_{z}=\nu^{eq}_{w,w^{\prime},z,u(z)}\right\}$
which carries an evaluation map. Using the evaluation map one can also define
@@
Note that ${\mathbb{Z}/p}\times T$ acts on $u:C\to X$ by
@@
and extended $q,t,\theta,\lambda$-linearly. The first sum is over a chosen basis of $H^{*}(X;\mathbb{F}_{p})$ as a vector space, which also generate $H^{*}_{T}(X;\mathbb{F}_{p})$ as a $H^{*}_{T}(\mathrm{pt};\mathbb{F}_{p})=\mathbb{F}_{p}[\![\lambda]\!]$-module by 2.12, implying equivariant formality.
The sign $\star$ is as in (2.22), but can be ignored under 2.12. We drop the superscript $T$ and abbreviate $\Sigma_{b}:=\Sigma_{b}^{T}$ in the following discussions when the group $T$ in concern is clear.
@@
For $b\in H^{*}(X;\mathbb{F}_{p})$, the operation $\Sigma_{b}^{T}$ satisfies the following properties.
@@
    $\Sigma_{1}^{T}=1$ for $1\in H^{0}(X;\mathbb{F}_{p})$, the unit class.
@@
    $\Sigma_{b}^{T}(b_{0})|_{q^{A}=0}=\mathrm{St}^{T}(b)\smile_{T}b_{0}$, where $\mathrm{St}^{T}(b)\in H^{*}_{T}(X;\mathbb{F}_{p}){[\![t,\theta]\!]}$ is the class obtained by the power operation construction as in (2.24) in $T$-equivariant cohomology.
@@
for each $a\in H^{2}(X;\mathbb{Z})$ defined by
@@
where $\ast_{T}$ is the $T$-equivariant quantum product, see 2.2.
@@ 2.4.2. Equivariant maps and shift operators
(2.48) $\mathcal{M}^{eq}(E_{\beta})=\left\{\left(u:C\to E_{\beta},\ (w,w^{\prime})\in E{\mathbb{Z}/p}\times ET\right):\pi_{E}(u(z))=z,\ \overline{\partial}_{\widetilde{J}_{w^{\prime}}}u=\nu_{E,w,w^{\prime}}^{eq}\right\}.$
For each homology class $\widetilde{A}\in H_{2}(E_{\beta};\mathbb{Z})$ of a section, so that $(\pi_{\beta})_{*}[\widetilde{A}]=[\mathbb{P}^{1}]$, there is a corresponding component of the moduli space where the degree of the map is prescribed by $u_{*}[C]=\widetilde{A}$. For the fixed cocharacter $\beta$, one can fix the minimal section $s_{\mathrm{min}}$ given by a fixed point $x_{\mathrm{min}}\in X^{\beta}$ where the moment map for $\rho\circ\beta$ attains a minimum.
@@
Denote the $({\mathbb{Z}/p}\times T)$-equivariant cohomology of $X$ with actions $\rho_{0}$ and $\rho_{\beta}$ by $H^{*}_{{\mathbb{Z}/p}\times T}(X|\rho_{0})$, $H^{*}_{{\mathbb{Z}/p}\times T}(X|\rho_{\beta})$ respectively. Note that by Künneth formula $H^{*}_{{\mathbb{Z}/p}\times T}(X|\rho_{0})\cong H^{*}_{T}(X|\rho){[\![t,\theta]\!]}$ since ${\mathbb{Z}/p}$ acts trivially on $X$ under $\rho_{0}$. Fix a cohomology class $b_{0}=\mathrm{PD}[Y_{0}]\in H^{*}_{{\mathbb{Z}/p}\times T}(X|\rho_{0})$ represented by a $T$-invariant locally finite cycle, and similarly fix $b_{\infty}=\mathrm{PD}[Y_{\infty}]\in H^{*}_{{\mathbb{Z}/p}\times T}(X|\rho_{\beta})$ represented by a $T$-invariant compact cycle. From the inclusion $\mathcal{Y}(E_{\beta})=Y_{0}\times Y_{\infty}\subseteq X_{0}\times X_{\infty}$ we construct an equivariant incidence cycle $\mathcal{Y}^{eq}(E_{\beta})$, cf. 2.3 and 2.16.
@@
The moduli space $\mathcal{M}^{eq;i;2j}_{A}(E_{\beta})\pitchfork\mathcal{Y}^{eq;i;2j}(E_{\beta})$ is regular of dimension $i+\sum_{k=1}^{r}2j_{k}+\dim_{\mathbb{R}}X+2c_{1}^{\mathrm{vert}}(s_{\mathrm{min}}+A)-|b_{0}|-|b_{\infty}|$, and can be compactified by stable maps and degeneration of equivariant parameters. In particular, when the dimension is 00, it gives a well-defined count in $\mathbb{F}_{p}$ (cf. 2.5).
@@
2.23 is a claim about compactness of the moduli spaces with incidence constraints, and will only hold for special choices of cocharacters $\beta:S^{1}\to T$ allowing maximum principle; see [Rit14, Section 1.12], [Iri17, Remark 3.10] for a discussion. For the example we are interested in computing, we will verify this compactness assumption in Section 3 for $\beta$ satisfying certain conditions on weights, see 3.16. In practice, 2.23 means that the shift operators can be defined without localization only for special choices of $\beta$. In general, defining the shift operators require passing to $H^{*}_{T}(X)_{\mathrm{loc}}$ for its definition. ∎
@@
(2.56) $\mathbb{S}_{\beta}\left(f(t,\lambda)\alpha\right)=f(t,\lambda+\beta t)\ \mathbb{S}_{\beta}\left(\alpha\right).$
In particular, it is not $H^{*}_{{\mathbb{Z}/p}\times{T}}(\mathrm{pt})$-linear.

### v3

@@ 2.2. Quantum Steenrod operations
Following [Lee23], we implement the incidence constraints via genuine cycles. We again fix a basis of $H^{*}(X;\mathbb{F}_{p})$ whose elements are denoted $b_{\bullet}$ for some subscript $\bullet$. Fix a basis cohomology class $b=\mathrm{PD}[Y]\in H^{*}(X;\mathbb{F}_{p})$ that is (a mod $p$ reduction of) Poincaré dual to a locally finite homology cycle $[Y]$ represented by a union $Y=\bigcup Y_{(i)}$ of embedded submanifolds $Y_{(i)}\subseteq X$ (using 2.2). Denote the map $Y\to X$ determined by the inclusions $Y_{(i)}\subseteq X$ as $Y\subseteq X$. Also fix $b_{0}\in H^{*}(X)$, $b_{\infty}\in H^{*}_{c}(X)$ Poincaré dual to a locally finite cycle $[Y_{0}]$ and (compact) cycle $[Y_{\infty}]$, respectively. Note that $b_{\infty}^{\vee}\in H^{*}(X)$ from the Poincaré pairing. Let $Y_{1}=\cdots=Y_{p}=Y$ be copies of the inclusion $Y\to X$.
@@
    For each $w\in E\mathbb{Z}/p$, $\mathcal{Y}_{w}$ is a product of maps $Y_{j}\to X$ representing a cycle homologous to the distinguished inclusion $Y_{j}\subseteq X$; denote the image of this product map also by $\mathcal{Y}_{w}$;
@@
Given an incidence cycle $\mathcal{Y}$, a corresponding equivariant incidence cycle $\mathcal{Y}^{eq}$ can be constructed inductively (in $i$) over the finite-dimensional cells $\Delta^{i}\subseteq E\mathbb{Z}/p$: we choose an arbitrary extension of the data defined over the $(i-1)$-skeleton of $E\mathbb{Z}/p$ to $\Delta^{i}\subseteq E\mathbb{Z}/p$ so that condition (ii) is satisfied, and equivariance condition (iii) fixes the data over all other $i$-cells in $E\mathbb{Z}/p$. Here we used $Y_{1}=\cdots=Y_{p}=Y$. In particular, $\mathcal{Y}^{eq}$ is only unique up to ($\mathbb{Z}/p$-equivariant) homotopy.
@@
(2.15) $\mathcal{M}^{eq}_{A}:=\left\{(u:C\to X,w\in E\mathbb{Z}/p):u_{*}[C]=A,\ (\overline{\partial}_{J}u)_{z}=\nu_{w,z,u(z)}\right\}.$
@@
    The evaluation map from moduli space of simple stable maps (with at least one bubble) is transverse to $\mathcal{Y}^{eq}$. In particular, when the counts $\mathcal{M}_{A}^{eq}\pitchfork\mathcal{Y}^{eq}$ are rigid, the simple stable maps only intersect $\mathcal{Y}^{eq}$ along $\mathcal{M}_{A}^{eq}$.
@@
Suppose 2.6 holds. Fix the unique $i\geq 0$ such that $\dim\mathcal{M}^{eq;i}_{A}\pitchfork\mathcal{Y}^{eq;i}=0$. Then the mod $p$ count of this 00-dimensional moduli space is independent of the choices of $J$, $\nu^{eq}$, and $\mathcal{Y}^{eq}$.
@@
    $\Sigma_{1}=1\cup=\mathrm{id}$ for $1\in H^{0}(X;\mathbb{F}_{p})$, the unit class.
@@
    $\Sigma_{b}(b_{0})|_{q^{A}=0}=\mathrm{St}(b)\cup b_{0}$, hence $\Sigma_{b}$ deforms the classical Steenrod operations in quantum parameters.
@@
The classical Steenrod operations in fact arise from the action of the symmetric group $S_{p}$ acting on $X^{p}$, not its subgroup $\mathbb{Z}/p\leq S_{p}$ of cyclic permutations. Since $H^{*}_{S_{p}}(\mathrm{pt};\mathbb{F}_{p})\cong\mathbb{F}_{p}[\![t^{p-1},\theta]\!]\hookrightarrow H^{*}_{\mathbb{Z}/p}(\mathrm{pt};\mathbb{F}_{p})\cong\mathbb{F}_{p}{[\![t,\theta]\!]}$, many coefficients of the expansion of $\mathrm{St}(b)$ from (2.24) in the equivariant parameters would in fact vanish. However, $\Sigma_{b}$ in general involves all monomials in $t,\theta$ that satisfy the dimension constraints (2.23). ∎
@@
(2.27) $\nabla_{a}\beta=t\partial_{a}\beta+a\ast\beta$
@@ 2.3. $T$-equivariant quantum Steenrod operations
(2.29) $D\rho\circ J_{\theta\cdot w^{\prime},x}=J_{w^{\prime},\rho(\theta)(x)}\circ D\rho:\ TX_{x}\to TX_{\theta\cdot x}.$
@@ A $(\mathbb{Z}/p\times T)$-equivariant inhomogeneous term is the data of
(2.30) $\nu^{eq}=\{\nu_{w,w^{\prime},z,x}\in\mathrm{Hom}^{0,1}(TC,TX)\}\in C^{\infty}((E\mathbb{Z}/p\times ET)\times C\times X;\mathrm{Hom}^{0,1}(TC,TX))$
@@
(2.31) $\nu_{\eta\cdot w,\theta\cdot w^{\prime},z,x}=D\rho(\theta)^{-1}\circ\nu_{w,w^{\prime},\tau(\eta)(z),\rho(\theta)(x)}\circ D\tau(\eta):TC_{z}\to TX_{x}.$
@@
Similarly, one can define the following analogue of the incidence cycles from 2.5. Fix a cohomology class $b=\mathrm{PD}[Y]\in H^{*}_{T}(X;\mathbb{F}_{p})$ that is Poincaré dual to a $T$-invariant locally finite homology cycle $[Y]$. Also fix $b_{0}\in H^{*}_{T}(X)$, $b_{\infty}\in H^{*}_{c,T}(X)$ Poincaré dual to a $T$-invariant locally finite cycle $[Y_{0}]$ and $T$-invariant cycle $[Y_{\infty}]$, respectively. Again fix the distinguished map $\mathcal{Y}=Y_{0}\times(Y\times\cdots\times Y)\times Y_{\infty}\subseteq X\times X^{p}\times X$; note that the image is $\mathbb{Z}/p\times T$-invariant.
@@
As the initial choice of $\mathcal{Y}$ is $(\mathbb{Z}/p\times T)$-invariant, we can construct the equivariant incidence cycle as a family of perturbations of $\mathcal{Y}$ parametrized over $E{\mathbb{Z}/p}\times ET$; see [Lee23, Lemma 5.5].
@@
(2.33) $\mathcal{M}^{eq}_{A}:=\left\{(u:C\to X,(w,w^{\prime})\in E\mathbb{Z}/p\times ET):u_{*}[C]=A,\ (\overline{\partial}_{J_{w^{\prime}}}u)_{z}=\nu_{w,w^{\prime},z,u(z)}\right\}$
which carries an evaluation map
(2.34) $\mathrm{ev}^{eq}:$ $\mathcal{M}_{A}^{eq}\to ET\times(X\times(E\mathbb{Z}/p\times X^{p})\times X)$
(2.35) $(u,w,w^{\prime})\mapsto(w^{\prime};u(z_{0});(w,u(z_{1}),\dots,u(z_{p}));u(z_{\infty})).$
Using the evaluation map one can also define
@@
Note that $(\eta,\theta)\in{\mathbb{Z}/p}\times T$ acts on $u:C\to X$ by
@@
and extended $q,t,\theta,\lambda$-linearly. The first sum is over a chosen basis of $H^{*}(X;\mathbb{F}_{p})$ as a vector space, which also generate $H^{*}_{T}(X;\mathbb{F}_{p})$ as a $H^{*}_{T}(\mathrm{pt};\mathbb{F}_{p})=\mathbb{F}_{p}[\![\lambda]\!]$-module by 2.2, which implies equivariant formality (2.3).
The sign $\star$ is as in (2.22), but can be ignored under 2.2. We drop the superscript $T$ and abbreviate $\Sigma_{b}:=\Sigma_{b}^{T}$ in the following discussions when the group $T$ in concern is clear.
@@
For $b\in H^{*}_{T}(X;\mathbb{F}_{p})$, the operation $\Sigma_{b}^{T}$ satisfies the following properties.
@@
    $\Sigma_{1}^{T}=\mathrm{id}$ for $1\in H^{0}_{T}(X;\mathbb{F}_{p})$, the unit class.
@@
    $\Sigma_{b}^{T}(b_{0})|_{q^{A}=0}=\mathrm{St}^{T}(b)\cup_{T}b_{0}$, where $\mathrm{St}^{T}(b)\in H^{*}_{T}(X;\mathbb{F}_{p}){[\![t,\theta]\!]}$ is the class obtained by the power operation construction as in (2.24) in $T$-equivariant cohomology.
@@
for each $a\in H^{2}_{T}(X;\mathbb{Z})$ defined by
@@
where $\ast_{T}$ is the $T$-equivariant quantum product, see 2.4. Here $\partial_{a}$ is the operator on $\Lambda$ which acts by $\partial_{a}q^{A}=\langle\bar{a},A\rangle q^{A}$ where $\bar{a}\in H^{2}(X)$ is the image of $a$ under $H^{2}_{T}(X)\to H^{2}(X)$.
@@ 2.4.2. Equivariant maps and shift operators
(2.50) $\mathcal{M}^{eq}(E_{\beta})=\left\{u:C\to E_{\beta},\ (w,w^{\prime})\in E{\mathbb{Z}/p}\times ET:\pi_{\beta}(u(z))=z,\ \overline{\partial}_{\widetilde{J}_{w^{\prime}}}u=\nu_{E,w,w^{\prime}}^{eq}\right\}.$
For each homology class $\widetilde{A}\in H_{2}(E_{\beta};\mathbb{Z})$ of a section (that is, $(\pi_{\beta})_{*}[\widetilde{A}]=[\mathbb{P}^{1}]$) there is a corresponding component of the moduli space where the degree of the map is prescribed by $u_{*}[C]=\widetilde{A}$. For the fixed cocharacter $\beta$, one can fix the minimal section $s_{\mathrm{min}}$ given by a fixed point $x_{\mathrm{min}}\in X^{T}\subseteq X^{\beta}$ where the moment map for $\rho\circ\beta$ attains a minimum over $X^{T}$ (which we assume to be compact).
@@
Denote the $({\mathbb{Z}/p}\times T)$-equivariant cohomology of $X$ with actions $\rho_{0}$ and $\rho_{\beta}$ by $H^{*}_{{\mathbb{Z}/p}\times T}(X|\rho_{0})$, $H^{*}_{{\mathbb{Z}/p}\times T}(X|\rho_{\beta})$ respectively. Note that by Künneth formula $H^{*}_{{\mathbb{Z}/p}\times T}(X|\rho_{0})\cong H^{*}_{T}(X|\rho){[\![t,\theta]\!]}$ since ${\mathbb{Z}/p}$ acts trivially on $X$ under $\rho_{0}$. Fix a cohomology class $b_{0}=\mathrm{PD}[Y_{0}]\in H^{*}_{{\mathbb{Z}/p}\times T}(X|\rho_{0})$ represented by a $T$-invariant locally finite cycle, and similarly fix $b_{\infty}=\mathrm{PD}[Y_{\infty}]\in H^{*}_{{\mathbb{Z}/p}\times T}(X|\rho_{\beta})$ represented by a $T$-invariant compact cycle. From the inclusion $\mathcal{Y}(E_{\beta})=Y_{0}\times Y_{\infty}\subseteq X_{0}\times X_{\infty}$ we construct an equivariant incidence cycle $\mathcal{Y}^{eq}(E_{\beta})$, cf. 2.5 and 2.16.
@@
The moduli space $\mathcal{M}^{eq;i;2j}_{A}(E_{\beta})\pitchfork\mathcal{Y}^{eq;i;2j}(E_{\beta})$ is regular of dimension $i+\sum_{k=1}^{r}2j_{k}+\dim_{\mathbb{R}}X+2c_{1}^{\mathrm{vert}}(s_{\mathrm{min}}+A)-|b_{0}|-|b_{\infty}|$, and can be compactified by stable maps and degeneration of equivariant parameters. In particular, when the dimension is 00, it gives a well-defined count in $\mathbb{F}_{p}$ (cf. 2.7).
@@
2.23 is a claim about compactness of the moduli spaces with incidence constraints, and will only hold for special choices of cocharacters $\beta:S^{1}\to T$ allowing maximum principle; see [Rit14, Section 1.12], [Iri17, Remark 3.10] for a discussion. For the example we are interested in computing, we will verify this compactness assumption in Section 3 for $\beta$ satisfying certain conditions on weights, see 3.19. In practice, 2.23 means that the shift operators can be defined without localization only for special choices of $\beta$. In general, defining the shift operators require passing to $H^{*}_{T}(X)_{\mathrm{loc}}$ for its definition. ∎
@@
(2.58) $\mathbb{S}_{\beta}\left(f(t,\lambda)\alpha\right)=f(t,\lambda+\beta t)\ \mathbb{S}_{\beta}\left(\alpha\right)$
for any polynomial $f$ in the variables $t,\lambda$. In particular, it is not $H^{*}_{{\mathbb{Z}/p}\times T}(\mathrm{pt})$-linear.